OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (1, -5)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

helpppppppppppppppppppppppppp

jimthompson5910 (jim_thompson5910):

if you were to plot (1,-5) and any point that is on the same horizontal level, say (3, -5), then you'd find that both points have the same y coordinate it turns out that all points on this horizontal line have a y coordinate of -5 so that means that the answer is y = -5

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

hey can u help me with 2 more

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

y < 1 over 3 x + 3 y less than or greater to 3 over 2 x + 3 y greater than or equal to 3 over 2 x + 3 The correct inequality is not listed.

jimthompson5910 (jim_thompson5910):

it says there's a cookie error and the picture won't show

OpenStudy (anonymous):

did u get it now

jimthompson5910 (jim_thompson5910):

ok that's better, thx

jimthompson5910 (jim_thompson5910):

choices b and c are contradictory....you can't have less and greater than at the same time

OpenStudy (anonymous):

ok can i ask u sum questions and u tell me if my answers are right

jimthompson5910 (jim_thompson5910):

sure, but some of those answer choices have a typo

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

http://learn.flvs.net/webdav/assessment_images/educator_algebra1/pictures/v8/04/00/04_00_group12_graph3.gifslope = -1 slope = 0 slope = undefined slope = 1

jimthompson5910 (jim_thompson5910):

getting another cookie error, i think i have to log in to see the image? that's probably what the error is saying, but idk

OpenStudy (anonymous):

oh ok ill jus write it out

jimthompson5910 (jim_thompson5910):

try taking a screenshot

OpenStudy (anonymous):

Which of the following is the slope between the points (0, -7) and (4, 12)? 19 fourths -19 fourths 4 nineteenths -4 nineteenths

OpenStudy (anonymous):

i got d

jimthompson5910 (jim_thompson5910):

use the slope formula to get m = (y2-y1)/(x2-x1) m = ( 12-( -7))/(4-0) m = ( 12+7 )/(4-0) m = 19/4

OpenStudy (anonymous):

]Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3) for this one i got d

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

ok Identify the correct slope and y intercept of the equation 6x - 3y = 12. slope = 2; y intercept at (0, -4) slope = -2; y intercept at (0, 4) slope = one half; y intercept at (0, 15) slope = one half; y intercept at (0, 9) i got c

jimthompson5910 (jim_thompson5910):

c is incorrect

OpenStudy (anonymous):

ok is it b

jimthompson5910 (jim_thompson5910):

get it into the form y = mx+b

jimthompson5910 (jim_thompson5910):

m = slope b = y-intercept

OpenStudy (anonymous):

ok so it will be a

jimthompson5910 (jim_thompson5910):

what equation did you get

jimthompson5910 (jim_thompson5910):

when you got it into the form y = mx+b

OpenStudy (anonymous):

ok i got 6x-3y=12 3 3 3 2x12

jimthompson5910 (jim_thompson5910):

no you first subtract 6x from both sides to get this -3y = -6x + 12 then you divide everything by -3 to fully isolate y

OpenStudy (anonymous):

ok i got 1y=2x+4

jimthompson5910 (jim_thompson5910):

12 divided by -3 is not 4

OpenStudy (anonymous):

oh sorry i meant negative 4

jimthompson5910 (jim_thompson5910):

so